# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v5

@@ 7 Hybrid Agency as Boundary-Condition Design
Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [khajehnejad2024]. Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency. In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components. Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input–output relation at the interface.
@@
The designer selects boundary conditions that reshape each component’s viable policy set, hoping that the intersection admits a collective policy that realises the desired input–output mapping at the interface.
@@ 8 Conclusions and Future Research
In this paper I used Stack Theory to describe MCAs as stacks of abstraction layers. In this view policies are embodied at each layer, approximating causality in service of adaptation, and cause is best approximated by weaker policies (w-maxing) [bennett2023c]. Formally, I prove a theorem (The Law of the Stack 5.1) showing w-maxing in higher layers requires sufficient w-maxing in lower layers. One reason artificial intelligence tends to be less adaptable is because it rests upon an inflexible stack of human abstractions. This claim is directed at contemporary, predominantly fixed-stack AI architectures. It is not a fundamental limitation on future computational or hybrid systems.
My claim is that, to build more robust systems, control should be delegated to the greatest extent possible. Obviously this must be balanced against desired functionality. Policies must be aligned. The system must still discriminate between correct and not. In the earlier sections I discussed the doctrine of mission command. What I am suggesting is that the concept of mission command should be applied to systems in general. At least, to the extent that adaptability is desired. Put another way, any constraints which are not strictly necessary to maintain the integrity of the system should be eliminated [^34] What counts as “strictly necessary” is relative to design goals and value judgements; see ESM Supplementary Note S4.. Humans tend to prefer ‘additive’ solutions [adams2021], by which I mean if a problem can be solved by adding a rule or subtracting a rule, we humans tend to add a rule. To effectively delegate control, we should subtract rules instead.
There are also software-first approaches that approximate aspects of this delegation picture, including federated learning [nori2025federated], constructivist AGI programmes [thorisson2012, goertzel2014], and active inference [Friston2010, friston2013]. An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6).
Future research might consider how adaptive hardware can be designed, to delegate control. Multilayer-causal-learning (MCL) provides a formal basis for ‘soft robotics’ [man2019]. Soft robotics is a proposed class of robots made of ‘soft’ materials that adapt. I have formalised adaptive materials as adaptation lower in the stack, providing a formal foundation for soft robotics. Self-assembling nanoparticle networks [borghi2024_brainlikehw] are a step in this direction. They can be used to form reconfigurable logic gates [paroli2023_receptron], facilitating adaptation at very low levels of abstraction. Taken to its logical conclusion, what might be required is a sort of artificial life, with small parts interacting form complex systems from the small scales and low levels of abstraction up [suzuki2006, ikegami2008, ikegami2007simulating].
As for the inductive bias mentioned in the introduction [chollet2019], it boils down to the choice of abstraction layer [bennett2024a]. I have described a means of learning a suitable abstraction layer. In theory, a multilayer-causal-learning system should be sufficient to solve ARC-AGI benchmark problems in terms of both sample and energy efficiency. Future research should test this empirically through (i) agent-based simulations of collectives with tunable delegation and abrupt environment shifts, measuring sample/energy efficiency and robustness; (ii) comparative benchmarks in which adaptation is confined to high-level software state versus systems with reconfigurable substrates (e.g. adaptive firmware or FPGAs); and (iii) hybrid experiments in which boundary conditions on biological components are systematically varied to quantify trade-offs. A more detailed set of candidate evaluation designs is provided in ESM Note S8.
Beyond hardware, cyberphysical systems in general can be evaluated in terms of abstraction, distribution and delegation. I might consider how states analogous to cancer could arise in human organisational structures, or how control can be delegated throughout an organisation to facilitate robustness and efficiency. For example, this could be applied to facilitate transformational bottom-up rather than incremental top-down adaptation to environmental changes.
Free-market trade is a great example of how delegation can facilitate adaptability in a large collective. As Hayek pointed out [hayek1945use], not all of the information is available to a central planner and so control is best delegated to the individual. The accumulation of resources within my formalism is implicit, rather than explicit. I have been concerned only with whether a system satisfies whatever arbitrary notion of correct I choose (i.e. does it achieve homeostatic and reproductive goals in whatever amount of time care about?), rather than how it evolves over time as a dynamical system. In other words, I have concerned myself only with whether a system arrives on time at the destination I want, not the journey it took to get there. Future research might formalise the accumulation of resources explicitly. One way to describe an evolving sequence of tasks over time could be something like a stack. To speculate about what my results imply for delegated systems like an economy, under less than ideal conditions resources may become concentrated in the hands of a few. The resulting system then delegates work, but not control. Those parts deprived of control may then not be able to construct a policy that aligns with the collective, and so they may splinter off and become ‘cancerous’. To prevent this, mechanisms can be put in place to preserve decentralisation without the need for intervention from outside the system (e.g. government intervening to break up a monopoly). Each system could be located along axes of redistribution and delegation, like an alternative political compass. Just enough redistribution of control and the system remains delegated. However, too much redistribution could have a similar effect, preventing parts from adapting to local conditions. An example of a mechanism that can effectively preserve delegation of control is adaptive rubber banding [mi2022]; a game mechanic (e.g. in Mario Kart) where success accrues penalties as well as resources, keeping the game competitive.
Another possibility is AI regulation. For example the UK government recently initiated “Safeguarded AI” [dalrymple2024b, dalrymple2024a], a project that aims to construct an AGI [goertzel2014] of level ASL-4 [anthropic2023] in a cyberphysical system, to safeguard against rogue AI. The idea of a “rogue AI” from within a cyberphysical system is analogous to cancer, adopting an “unsafe” policy at odds with that of the collective. AI regulations may over-constrain systems and trigger the very cancerous state they seek to prevent. To reiterate the earlier point about subtractive solutions [adams2021], any guardrails which are not strictly necessary must be eliminated to make room for those that are.
@@
In summary, I have described systems in terms of delegation and abstraction. I proved that adaptability at higher levels of abstraction is constrained by adaptability at lower levels (The Law of the Stack). I have argued that one reason biological systems are more ‘intelligent’, meaning sample and energy efficient, is because they delegate adaptation down the stack. This Stack Theoretic perspective lays a foundation for future empirical research, showing how we might build more robust systems of any kind.
9 Electronic Supplementary Materials
9.1 Mathematical Results
The Law of the Stack
Theorem 9.1 (The Law of the Stack).
Assume the formal setting defined in the main text: finite vocabulary $v\subseteq P$, language $L_{v}=\{\,\ell\subseteq v:\bigcap\ell\neq\emptyset\,\}$, extensions $E_{\pi}=\{\,y\in L_{v}:\pi\subseteq y\,\}$, $v$-tasks $\gamma=\langle I_{\gamma},O_{\gamma}\rangle$, correct policies $\Pi_{\gamma}=\{\pi\in L_{v}:E_{I_{\gamma}}\cap E_{\pi}=O_{\gamma}\}$, and utility $\epsilon(\gamma)=\max_{\pi\in\Pi_{\gamma}}(|E_{\pi}|-|O_{\gamma}|)$. Let $\langle\lambda_{0},\dots,\lambda_{n}\rangle$ be an MLA with abstractor $\mathfrak{f}$ and state $\langle(\pi_{0},\dots,\pi_{n}),(v_{0},\dots,v_{n})\rangle$ such that $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$ and $\pi_{i}\in\Pi_{\lambda_{i}(v_{i})}$ for all $i<n$.
Fix any $i<n$, and define $\gamma_{i+1}=\lambda_{i+1}(v_{i+1})$ where $v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})$. Then
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\;\leq\;2^{|E_{\pi_{i}}|}.$
Equivalently,
$|E_{\pi_{i}}|\;\geq\;\log_{2}\!\big(\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\big).$
Thus, higher utility at layer i+1i\!+\!1 requires (at least) a weaker policy at layer $i$ in the sense of larger extension cardinality.
Proof.
We proceed by a sequence of elementary bounds that make the layer-to-layer dependency explicit.
Lemma 1 (Languages are bounded by the power set).
For any finite vocabulary $v$, the language $L_{v}$ satisfies $|L_{v}|\leq 2^{|v|}$.
Proof. By definition, $L_{v}\subseteq\mathcal{P}(v)$ because every $\ell\in L_{v}$ is a subset of $v$. Since $|\mathcal{P}(v)|=2^{|v|}$ for finite $v$, we have $|L_{v}|\leq 2^{|v|}$. $\square$
Lemma 2 (Extensions live inside the language).
If $\pi\in L_{v}$, then $E_{\pi}\subseteq L_{v}$, hence $|E_{\pi}|\leq|L_{v}|$.
Proof. By definition $E_{\pi}=\{y\in L_{v}:\pi\subseteq y\}$, so $E_{\pi}\subseteq L_{v}$. Taking cardinalities yields $|E_{\pi}|\leq|L_{v}|$. $\square$
Lemma 3 (Utility implies a lower bound on language size).
Let $\gamma$ be a $v$-task. Then
$\epsilon(\gamma)+|O_{\gamma}|\leq|L_{v}|.$
Proof. Because $v$ is finite, $L_{v}$ is finite, hence $\Pi_{\gamma}\subseteq L_{v}$ is finite, so the maximum defining $\epsilon(\gamma)$ is attained by some $\pi^{\star}\in\Pi_{\gamma}$. By definition,
$\epsilon(\gamma)=|E_{\pi^{\star}}|-|O_{\gamma}|\quad\Rightarrow\quad\epsilon(\gamma)+|O_{\gamma}|=|E_{\pi^{\star}}|.$
By Lemma 2, $|E_{\pi^{\star}}|\leq|L_{v}|$. Combining yields the claim. $\square$
Lemma 4 (The abstractor output vocabulary is an image of an extension).
Let $T:L_{v_{i}}\to P$ be the map $T(o)=\bigcap o$. Under the MLA definition,
$v_{i+1}=\mathfrak{f}(v_{i},\pi_{i})=\{\,T(o):o\in E_{\pi_{i}}\,\},$
so $v_{i+1}$ is the image of $E_{\pi_{i}}$ under $T$. Consequently,
$|v_{i+1}|\leq|E_{\pi_{i}}|.$
Proof. This is just re-writing the definition of $\mathfrak{f}$ in functional form, and then applying the general fact that for finite sets $A$ and any function $g$, $|g(A)|\leq|A|$. $\square$
Conclusion.
Apply Lemma 3 to $\gamma_{i+1}$ in vocabulary $v_{i+1}$ to obtain
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq|L_{v_{i+1}}|.$
Apply Lemma 1 to bound $|L_{v_{i+1}}|\leq 2^{|v_{i+1}|}$, giving
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|v_{i+1}|}.$
Finally apply Lemma 4: $|v_{i+1}|\leq|E_{\pi_{i}}|$. Since $2^{x}$ is monotone increasing,
$2^{|v_{i+1}|}\leq 2^{|E_{\pi_{i}}|}.$
Combining the last two displays yields
$\epsilon(\gamma_{i+1})+|O_{\gamma_{i+1}}|\leq 2^{|E_{\pi_{i}}|},$
and the logarithmic form follows by applying $\log_{2}$ (also monotone). ∎
Remark. In the thesis version of this result, the intuition is phrased as: larger vocabularies at layer i+1i{+}1 permit weaker (more general) policies, hence more adaptation at higher layers. The proof here makes the dependency explicit by bounding the achievable utility at layer i+1i{+}1 as a monotone function of the extension size $|E_{\pi_{i}}|$ at layer $i$, without requiring any additional monotonicity assumptions about $\epsilon(\lambda(\cdot))$ under set inclusion.
Hybrid case study formalisation (Pong organoid)
Let $C$ be a set of cells. For each cell $i\in C$, let $\Pi_{i}$ denote the set of viable policies for that cell under some interface/environmental conditions. The set of viable collective policies for the organoid is $\Pi_{C}=\bigcap_{i\in C}\Pi_{i}$. A conventional computer interface (electrodes, stimulus generator, game process) imposes boundary conditions $c$ that restrict which policies remain viable; write this restriction schematically as an operator $R_{c}$ such that $\Pi_{i}^{(c)}=R_{c}(\Pi_{i})$ and $\Pi_{C}^{(c)}=\bigcap_{i\in C}\Pi_{i}^{(c)}$. The design objective is to choose $c$ so that (i) $\Pi_{C}^{(c)}\neq\emptyset$ and (ii) the remaining collective policies implement a target output set $O_{\text{pong}}$ at the interface. In this view, building the hybrid is a constraint-design problem: impose the weakest (least exclusionary) boundary conditions that make $O_{\text{pong}}$ achievable while preserving non-empty intersection (collective identity).
9.2 Supplementary Notes and Extended Discussion
@@
There are many definitions of intelligence, and the merits of each are beyond the scope of this paper. Mine is based on [bennett2025thesis], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation. It is similar to Wang’s definition of intelligence as adaptation with limited resources [wang2019], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance. By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems. This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
@@
The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [bennett2025thesis]. A higher level of abstraction is the behaviour of the layers below. For example, an organ like a heart is a behaviour of cells, human is a behaviour of some organs, and language is a behaviour of humans.
@@
Weakness is formally defined in the appendix. Related work proved w-maxing to be optimal regardless of the abstraction layer [bennett2025e, bennett2023b, bennett2024a, bennett2025thesis]. In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness. If $x$ is weaker than $y$, that means it commits to less, but achieves the same end. For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [bennett2023b]). This is a bit like the principle of maximum entropy [jaynes1957a, jaynes1957b], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [bennett2024b].
For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [bennett2024appendices]. It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
@@
As Hume argued, a description of what ought to be cannot be derived from only descriptions of what is. “Strictly necessary” is a value judgement, or an implicit choice of what ought to be. In the thesis [bennett2025thesis] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment. These transitions preserve some aspects of that environment, while destroying others. Here however I am more concerned with the design of systems which serve a human purpose. Strictly necessary is a matter of desired functionality.
@@
Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [bennett2025thesis]. The full justification for these definitions is out of scope and addressed at length elsewhere [bennett2024a, bennett2025thesis, bennett2025e]. However I will provide a brief overview. To make claims that hold regardless of the abstraction layer, we must rely only on what holds for every abstraction layer. From our subjective perspective we can’t know what is at the bottom of our stack (it may be infinite). I call this bottom layer the environment. Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction. Hence there must be a set of states where each state is one point of difference [^35] Nothing is assumed about what these states might ontologically be or epistemologically contain.. Because these states must be mutually exclusive, this frames time as difference. Each possible sequence of states is a timeline or possible world. Truth is defined ontologically, relative to states. A set of states is effectively a set of truth conditions, so a possible fact is called a declarative program, and a program is true given a state $\phi$ if it contains that state. Every aspect of the environment is just a set of programs. Everything that exists is a set of facts, and each timeline preserves different sets of facts. A timeline is like an opinion; as time progresses some aspects of the environment are preserved and others are destroyed. Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [bennett2025thesis]. If the set of states is infinite then there are infinite possible worlds. Assume I have a robotic system. I get an abstraction layer for the system by choosing a finite aspect $\mathfrak{v}$ of programs called a vocabulary. Everything the abstraction layer can express is a subset of its vocabulary. For example if a system has an arm it can raise or lower, then there are programs in $\mathfrak{v}$ which correspond to actuation of the arm in some form, though it need not be neat symbolic representations like ’raise arm’ or ’lower arm’. Since states are mutually exclusive, $\mathfrak{v}$ gives us an embodied formal language. I do not mean like the words humans use to communicate. I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [nakagaki2000] [^36] Using this frame cognition can be formalised in terms that align with both enactivist [thompson2007] and pancomputationalist [piccinini2021] perspectives. Hence, cognition within Stack Theory is also called Pancomputational Enactivism[bennett2024a].
@@
Federated machine learning delegates control to some extent (e.g. shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [nori2025federated]. Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [thorisson2012, goertzel2014]. The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [nivel2013] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [thorisson2018cumlearncause], “growing” knowledge in a collective meta-model through cumulative continuous learning [thorisson2019cumlearning], and explicitly targeting adaptability [thorisson2014autonomousNatCom, thorisson2020, eberding2024, sheikhlar2024] [^37] See also http://www.openaera.org.. ActPC-Chem [goertzel2024] is a recent theory behind the distributed Hyperon architecture [goertzel2023] that was influenced by theoretical and methodological insights from AERA.
@@
Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [Friston2010, friston2013, parr2024]. The present paper addresses a different axis: it states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1). In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer. An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers; conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
@@
Several lines of work suggest points of contact, without implying formal equivalence. ActPC-Chem frames delegation in explicitly Active Inference terms [goertzel2024, Friston2010, friston2013]. “Scale-free” active inference proposes that similar dynamical principles repeat across scales [friston2023b, friston2024a], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity. Delegated [ororbia2022, ororbia2023, ororbia2023b] or federated [friston2024b] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale. Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g. the Landauer bound) by reducing memory–processor separation and associated bottlenecks [ororbia2024mortalcomputationfoundationbiomimetic, aurell2012, zou2021], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead. However, the theories still differ in their primitives (e.g. tasks/weakness versus generative models/free energy), and it remains an open question whether weakness (w-maxing) can be recovered as, or justified by, a principle within the Active Inference apparatus. Establishing such a mapping is an important direction for future work.

### v6

@@ 7 Hybrid Creation as Weak Boundary-Condition Design
Consider the aforementioned example of the organoid adapted to play the game Pong with notable sample efficiency relative to conventional reinforcement-learning baselines [53]. Hybrid systems like organoids are not fixed interpreters of software-level policy, but collectives whose lower-level parts retain bottom-up agency. In such systems, hybrid creation is the act of designing an interface and environment that reshapes the feasible policy space of the biological components. Rather than “programming” the parts directly, the designer engineers boundary conditions so that the remaining viable collective policies implement a target input-output relation at the interface. Practically, hybrid creation is iterative:
  1. 1.
    choose boundary conditions, then
  2. 2.
    observe emergent collective behaviour, then
  3. 3.
    tighten or relax constraints to steer the intersection of viable policy sets toward the target interface behaviour and
  4. 4.
    repeat.
Similar boundary-condition design issues arise in other bio-hybrid constructs (e.g. xenobot-like cell collectives) and in soft-robotic substrates where not just software but morphology and material properties shape the feasible low-level policy space.
@@
The designer selects boundary conditions that reshape each component’s viable policy set, hoping that the intersection admits a collective policy that realises the desired input-output mapping at the interface.
@@ 8 Conclusions and Future Research
Biological systems often adapt efficiently because they delegate adaptive control across scales. Using Stack Theory [19], I formalised delegation as the depth to which adaptation is permitted in a stack of abstraction layers and proved The Law of the Stack (Theorem 5.1). Utility one layer up is bounded by the weakness of the realised policy one layer down. Delegation is a structural prerequisite for high-level adaptability.
This constraint is compatible with active inference. Under a simple viability prior, minimising variational free energy is equivalent to maximising entropy over viable futures (w-maxing; ESM “Viability, lived history, and variational free energy”). Corollary 5.1 shows that free energy minimisation inherits the same bottleneck: abstract-layer free energy cannot be reduced without sufficient lower-layer slack.
The implication for AI is not that computation cannot be intelligent, but that contemporary AI stacks tend to concentrate adaptation at high software layers while freezing lower layers, thereby throttling achievable weakness and adaptation efficiency. For hybrid agents (organoids, soft robots, and human–AI organisations) the design problem becomes boundary-condition engineering: constrain parts just enough to elicit target collective behaviour while preserving a non-empty intersection of viable policies (collective identity). When that intersection collapses, the system can fragment into competing subagents (Section 6).
Future work should test these predictions by varying which layers are permitted to adapt (e.g. hardware-in-the-loop learning, morphologically adaptive substrates, and biologically integrated controllers) and measuring resulting degrees of freedom, free energy, and robustness. More broadly, this paper is a stepping stone toward a unified account of delegation, weakness/max-entropy heuristics, and multiscale agency in the larger Stack Theory programme [19].
A practical design implication is that, when adaptability is desired, constraints should be weakened as far as possible without breaking the system’s intended function. Constraints that are not strictly necessary relative to the design goal should be removed (What counts as “strictly necessary” is relative to design goals and value judgements. See ESM Supplementary Note S4.). Humans often prefer ‘additive’ solutions [1], meaning if a problem can be solved by adding a rule or subtracting a rule, we tend to add a rule. However The Law of the Stack suggests a preference for subtractive solutions might be beneficial in many circumstances.
To speculate about future research beyond experimental verification, Stack Theory is a lens that can be applied generally. For example, a number of software-first approaches to AI already approximate aspects of this delegation picture, including federated learning [52], constructivist AGI programmes [90, 42], and active inference [39, 40]. An expanded discussion (including AERA, Hyperon/ActPC-Chem, and links to scale-free active inference) is provided in the ESM (Supplementary Note S6). Stack Theory might also inform the design of ‘soft robots’ [58]. Soft robots are a proposed class of robots made of ‘soft’ materials that adapt. I have formalised adaptive materials as adaptation lower in the stack, which can be used as a formal foundation for soft robotics. Self-assembling nanoparticle networks [24] are a step in this direction. They can be used to form reconfigurable logic gates [70], facilitating adaptation at very low levels of abstraction. Taken to its logical conclusion, future AI may more closely resemble artificial life [85, 45, 46]. Finally, Stack Theory may be extended to economic and organisational systems. For example, free-market trade could be seen as a classic argument for delegation. As Hayek pointed out [44], a great deal of relevant information is spread among a population, unavailable to a central planner. This motivates delegation of control, to shorten feedback loops. However, the present formalism treats resources implicitly via task correctness, so connecting these arguments to economic dynamics would require extending Stack Theory to explicit time-series [23] and resource accumulation. Additional speculative extensions and applications (including organisational and economic systems, AI safety regulation, and temporal extensions of stacks) are collected in ESM Note S9.
@@
In summary, I have applied Stack Theory to explore how delegation of control affects adaptability. The Law of the Stack shows how adaptability at higher levels of abstraction is constrained by adaptability at lower levels, and delegation is structurally necessary to minimise free energy. This lays a foundation for a broad range of future theoretical and empirical research, showing how systems of any kind might be made more robust and adaptable.
Acknowledgements
First, I’d like to thank the anonymous peer reviewers for their detailed and helpful feedback. Second, I’d like to thank Anna Ciaunica, Peter Watts, Sean Welsh, Ricard Solé, Karl Friston, Antonio Damasio, Kingson Man, Paolo Milani, Ben Goertzel, Kristinn R. Thórisson, Elija Perrier and Noel Hinton for their support and feedback on this paper.
@@ Acknowledgements
There are many definitions of intelligence, and the merits of each are beyond the scope of this paper. Mine is based on [19], which frames intelligence as the ability to complete a wide range of tasks, which is sample and energy efficiency in adaptation. It is similar to Wang’s definition of intelligence as adaptation with limited resources [96], and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. Mainstream AI often operationalizes ‘intelligence’ as general-purpose task performance. By that standard, one may argue that current large language models are ‘more intelligent’ than many biological systems. However that definition is fundamentally flawed, as it fails to express the comparative efficiency with which humans acquire general purpose competence [16, 96, 26]. In theory even a lookup table or rote-learning response can possess general purpose competence, given enough resources. This paper instead treats intelligence as the cost of adaptation (sample and energy efficiency), and the results should be read accordingly.
@@
The notion of levels of abstraction is more typically used in pure computer science settings, but can be applied to any system [19]. A higher level of abstraction is the behaviour of the layers below. For example, an organ like a heart is a behaviour of cells, human is a behaviour of some organs, and language is a behaviour of humans.
@@
Weakness is formally defined in the appendix. Related work proved w-maxing to be optimal regardless of the abstraction layer [18, 15, 16, 19]. In narrative terms, assume two policies $x$ and $y$ both constrain the system to examples of “correct” behaviour given past inputs, but differ in weakness. If $x$ is weaker than $y$, that means it commits to less, but achieves the same end. For example, if I assume the lawyer was paid at least $\mathdollar 5$, I am more weakly constrained than if I assume that lawyer was paid at least $\mathdollar 6$.
This example also helps separate weakness from simplicity. The difference between $\mathdollar 5$ and $\mathdollar 6$ is not primarily about how hard the rule is to write down. It is about how many further commitments remain compatible with the rule, given the agent’s available distinctions. Weakness is defined relative to an abstraction layer. The layer fixes what counts as a meaningful refinement because it fixes what the system can jointly discriminate and jointly realise. If the embodiment changes, then the induced vocabulary changes, and so does the completion structure that weakness counts.
Many familiar heuristics can be seen as special cases of the same idea once this completion structure is made explicit. If one treats all completions that remain possible at a layer as equally plausible, then choosing a correct policy that maximises $|\mathrm{Ext}(\pi)|$ is equivalent to maximising entropy over viable futures inside that layer. This parallels least commitment and version space intuitions, but it is stated directly in terms of the agent’s embodied language rather than in terms of an encoding chosen by an observer. That is what lets weakness be used across layers in Stack Theory, where a policy at one layer induces the next layer vocabulary via the abstractor map.
In the language of agency, weakness can be read as a quantitative proxy for autonomy at a layer. A weaker policy leaves more degrees of freedom (more “ways of being right”) while still satisfying the same observed constraints. In multiscale and hybrid systems this autonomy matters because it increases the chance that component policies remain compatible with a higher-level collective policy, preserving collective identity.
As an epistemological razor, this means “explanations should be no more specific than necessary” or alternatively “rules should constrain no more than necessary” (see Bennett’s Razor in [15]). This is a bit like the principle of maximum entropy [50, 49], difference being that w-maxing also accounts for the abstraction layer, which is important because abstraction can render complexity subjective [17].
For a rigorous definition of learning and efficiency, see definitions in the online repository for the thesis [21]. It includes proofs showing w-maxing is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
@@
As Hume argued, a description of what ought to be cannot be derived from only descriptions of what is. “Strictly necessary” is a value judgement, or an implicit choice of what ought to be. In the thesis [19] on which this is based this ought from which other descriptions of ought follow is derived from the state transitions of the environment. These transitions preserve some aspects of that environment, while destroying others. Here however I am more concerned with the design of systems which serve a human purpose. Strictly necessary is a matter of desired functionality.
@@
Stack Theory frames the description of an objective reality in terms of choices of subjective lens (an abstraction layer) [19]. The full justification for these definitions is out of scope and addressed at length elsewhere [16, 19, 18]. However I will provide a brief overview. To make claims that hold regardless of the abstraction layer, we must rely only on what holds for every abstraction layer. From our subjective perspective we can’t know what is at the bottom of our stack (it may be infinite). I call this bottom layer the environment. Whatever the environment is, there must be change or difference for there to be anything at higher levels of abstraction. Hence there must be a set of states where each state is one point of difference (Nothing is assumed about what these states might ontologically be or epistemologically contain.). Because these states must be mutually exclusive, this frames time as difference. Each possible sequence of states is a timeline or possible world. Truth is defined ontologically, relative to states. A set of states is effectively a set of truth conditions, so a possible fact is called a declarative program, and a program is true given a state $\phi$ if it contains that state. Every aspect of the environment is just a set of programs. Everything that exists is a set of facts, and each timeline preserves different sets of facts. A timeline is like an opinion. As time progresses some aspects of the environment are preserved and others are destroyed. Each timeline preserves different aspects, expressing a different “ought” which related work uses to explain the origins of goal directed behaviour [19]. If the set of states is infinite then there are infinite possible worlds. Assume I have a robotic system. I get an abstraction layer for the system by choosing a finite aspect $\mathfrak{v}$ of programs called a vocabulary. Everything the abstraction layer can express is a subset of its vocabulary. For example if a system has an arm it can raise or lower, then there are programs in $\mathfrak{v}$ which correspond to actuation of the arm in some form, though it need not be neat symbolic representations like ’raise arm’ or ’lower arm’. Since states are mutually exclusive, $\mathfrak{v}$ gives us an embodied formal language. I do not mean like the words humans use to communicate. I mean a language of embodiment, where every statement in that language constrains rather than describes the system, like the aforementioned boundary conditions on slime mold [62] (Using this frame cognition can be formalised in terms that align with both enactivist [86] and pancomputationalist [73] perspectives. Hence, cognition within Stack Theory is also called Pancomputational Enactivism[16].)
@@
Federated machine learning delegates control to some extent (e.g. shifting data ownership and parts of optimisation to edge devices), which can improve privacy and efficiency [52]. Constructivist theories of artificial general intelligence emphasise multilayered self-organisation at the level of software while acknowledging at least some of the constraints imposed by hardware [90, 42]. The Auto-catalytic Endogenous Reflective Architecture (AERA) is a goal-driven ampliative reasoning system [64] based on constructivist principles that relies on causal knowledge representation to achieve autonomy [89], “growing” knowledge in a collective meta-model through cumulative continuous learning [88], and explicitly targeting adaptability [87, 91, 33, 80] (See also http://www.openaera.org.). ActPC-Chem [43] is a recent theory behind the distributed Hyperon architecture [41] that was influenced by theoretical and methodological insights from AERA.
@@
Active Inference (and the Free Energy Principle, FEP) provides a normative account of perception and action in which an agent maintains a generative model and selects actions to minimise variational free energy (and, by extension, expected free energy) [39, 40, 71]. The present paper addresses a different axis. It states a structural constraint on stacks of abstraction layers, namely that what can be achieved at layer i+1i{+}1 (utility $\epsilon(\gamma_{i+1})$) is bounded by the extension (weakness) of the realised policy at layer $i$ (Theorem 5.1). In this sense, MCL is about where adaptation is implemented (and how much latitude lower layers retain), rather than about the particular inferential objective used within a layer. An Active Inference agent implemented on a predominantly fixed computational substrate can still be bottlenecked by static lower layers. Conversely, a delegated multiscale substrate can in principle host a range of learning rules, including Active Inference.
@@
Several lines of work suggest points of contact, without implying formal equivalence. ActPC-Chem frames delegation in explicitly Active Inference terms [43, 39, 40]. “Scale-free” active inference proposes that similar dynamical principles repeat across scales [36, 37], which is conceptually close to the MCA picture and to the idea that policy alignment must be maintained across layers to preserve collective identity. Delegated [68, 69, 67] or federated [38] Active Inference approaches may therefore be promising candidates for building efficient hybrid systems at scale. Separately, work on “mortal computation” seeks to approach thermodynamic limits (e.g. the Landauer bound) by reducing memory and processor separation and associated bottlenecks [66, 5, 97], which resonates with the present paper’s emphasis on efficiency and with the intuition that pushing adaptation downward can reduce overhead. However, the theories still differ in their primitives (e.g. tasks and weakness versus generative models and free energy). Even so, a simple mapping is available when viability is treated as a hard constraint and preferences are uniform across viable futures, see the Mathematical Results subsection “Viability, lived history, and variational free energy”. A deeper identification in the general non-uniform case remains open.